Assume that a single Inhibitory Postsynaptic Potential (IPSP) has a negative magnitude of -0.5mV at the axon hillock, and that a single Excitatory Postsynaptic Potential (EPSP) has a positive magnitude of +0.5mV. For a neuron with initial membrane potential of -70mV, calculate the net effect of the simultaneous arrival of 6 IPSPs and 2 EPSPs on the membrane potential. Show your work.
Answer #1

Strength of a single IPSP= -0.5mV

Hence, strength of 6 IPSPs= 6*(-0.5) = -3.0mV

Strength of a single EPSP= +0.5mV

Hence, strength of 2 EPSPs= 2*(+0.5) = +1.0mV

Now, combined strength of signals

= -3.0mV +1.0mV = -2.0mV

Therefore, for a neuron with initial membrane potential of -70mV, the arrival of the signals would result in a final membrane potential of -72mV.